OpenStudy (anonymous):

solve solution 3x+3y=27 x-3y=-11

OpenStudy (dtan5457):

There are many methods to do this. Ever heard of substitution or elimination ?

jimthompson5910 (jim_thompson5910):

Add the equations to get 3x+3y=27 x-3y=-11 --------- 4x+0y = 16 so we know that 4x = 16 ---> x = ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

x = 4 is correct

jimthompson5910 (jim_thompson5910):

now you use this x value to find the y value

jimthompson5910 (jim_thompson5910):

3x+3y=27 3(4)+3y=27 ... replace x with 4 12 + 3y = 27 y = ??

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

I replaced x with 4 in the first equation 3x+3y=27

jimthompson5910 (jim_thompson5910):

the goal from there is to solve for y

OpenStudy (anonymous):

is it 4

jimthompson5910 (jim_thompson5910):

12 + 3y = 27 12 + 3y-12 = 27-12 ... Subtract X from both sides. 3y = 15 y = ??

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

yes, together x = 4 and y = 5 the solution is the ordered pair (4,5)

jimthompson5910 (jim_thompson5910):

oops typo, meant to say "Subtract 12 from both sides." but you get what I mean hopefully